OpenStudy (anonymous):

Find cos theta if sin theta = -12/13 and tan theta > 0

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

Hello, how would I solve this? Should I be drawing a triangle?

OpenStudy (misty1212):



OpenStudy (anonymous):

why is 13 on the long side? Is it not the top of the fraction going to the longer side?

OpenStudy (jhannybean):

missed the negative?

OpenStudy (campbell_st):

you need to recognise this this is a 3rd quadrant angle and a iwll be negative... and 12 should be shown as a -12

OpenStudy (misty1212):

sine is opposite over hypotenuse forget the negative, find the length of the other side you can decide if your answer is positive or negative once you get the length of the other leg

OpenStudy (anonymous):

wait, sin = opposite over hypothesis so that is why 13 is on the long side, correct?

OpenStudy (misty1212):

yes

OpenStudy (misty1212):

"long side" you mean "hypotenuse" right?

OpenStudy (anonymous):

i would use Pythagorean right? so

OpenStudy (jhannybean):

And that only applies to right angle triangle.

OpenStudy (anonymous):

\[\sqrt{13}^2 - \sqrt{12}^2\]

OpenStudy (misty1212):

yes or recall this very famous right triangle not as famous as 3- 4 - 5 but almost

OpenStudy (misty1212):

no \[\sqrt{13^2-12^2}\]

OpenStudy (anonymous):

yes, i tried to write that but i didn't know how, sorry

OpenStudy (misty1212):

\[a^2+12^2=13^2\\ a^2=13^2-12^2\\ a=\sqrt{13^2-12^2}\]

OpenStudy (misty1212):

you will see this nice triangle again and again almost as often as you see \(3:4:5\) and it relatives \(6:8:10\) etc

OpenStudy (anonymous):

a = 5

OpenStudy (anonymous):

cos = 5/13?

OpenStudy (misty1212):



OpenStudy (anonymous):

but it is negative right? why?

OpenStudy (misty1212):

no be careful yes it is negative

OpenStudy (misty1212):

it is negative because you are told that tangent is positive, and tangent is sine over cosine, so if tangent is positive and sine is negative, then cosine must be negative too

OpenStudy (anonymous):

really? that makes sense, thank you very much, i think i understand how to do this now, for any problem i would repeat similar steps

OpenStudy (misty1212):

or if you like to think of it this way, if sine is negative and tangent is positive, you must be in quadrant 3 it is the same thing really you are welcome

OpenStudy (anonymous):

i have one quick question

OpenStudy (anonymous):

can you answer it quickly?

OpenStudy (misty1212):

if i know it

OpenStudy (anonymous):

for a problem like \[\frac{ \tan \Theta }{ \sec \Theta } \]

OpenStudy (anonymous):

i would need to know the fundamentals right? or would i be able to solve it without it?

OpenStudy (misty1212):

to simplify it is usually algebra

OpenStudy (misty1212):

you need to know how to write each of these in terms of sine and cosine, most of the problems require little or no trig, only algebra

OpenStudy (misty1212):

for example you have \[\huge\frac{\frac{\sin(x)}{\cos(x)}}{\frac{1}{\cos(x)}}\]

OpenStudy (anonymous):

sin(x)/1

OpenStudy (misty1212):

the cosines cancel and you get sin(x)

OpenStudy (anonymous):

but how do you know sec = 1/cos(x)

OpenStudy (misty1212):

if you write sin(x)/1 your teacher will think you are kind of dumb, so don't do it

OpenStudy (anonymous):

haha you're right

OpenStudy (misty1212):

it is the definition of secand

OpenStudy (misty1212):

secant

OpenStudy (misty1212):

that is why you do not have a secant button on your calculator, it is not necessary it is the reciprocal of cosine

OpenStudy (anonymous):

So if i googled the definitions of sec, sine, cos, and more i could solve these questions?

OpenStudy (jhannybean):

there are a whole lost of trig ratios and identities you can find on gogle.

OpenStudy (jhannybean):

Google*

OpenStudy (misty1212):

you need to memorize them

OpenStudy (anonymous):

okay, thank you very much again

OpenStudy (misty1212):

you are welcome again good luck with your trig class

OpenStudy (anonymous):

wow wow wow! Thank you for the sheet!!!

OpenStudy (misty1212):

make sure to look at the unit circle on the last page

OpenStudy (misty1212):

that gives you the values for sine and cosine for lots of angles, both in degrees and in radians

OpenStudy (anonymous):

Thank you very very much! Merry almost Christmas and happy Hanuka!

OpenStudy (misty1212):

merry christmas to you too or whatever

OpenStudy (anonymous):

lol